OpenStudy (anonymous):

-12 + m = - 87 A. m = - 75 B. m = - 99 C. m = 75 D. m = 99

OpenStudy (anonymous):

-12 + m = -87 +12 +12 m = - 75

OpenStudy (anonymous):

-12 + m = - 87 +12 +12 m=-75

OpenStudy (anonymous):

ha u beat me

OpenStudy (anonymous):

haha :) yes i did